OpenStudy (mendicant_bias):

Integration problem regarding substitution. Posting below in a minute or two:

OpenStudy (mendicant_bias):

"Use the substitution formula in Theorem 7 to evaluate the integral." (Substitution formula is \[\int\limits_{a}^{b}f(g(x))*g'(x)dx = \int\limits_{g(a)}^{g(b)}f(u)du.\] Where the integral given is:\[\int\limits_{-1}^{0}\sqrt{y+1} dy\]I'm not sure what i'm doing wrong, but here's how i' m trying to do it: \[f(x) = \sqrt{u}\]\[u = (y+1)\] \[g(-1) = (-1)+1 = 0 \]\[g(0) = (0)+1 = 1\] \[\int\limits_{0}^{1}u ^{(1/2)}du = \int\limits_{0}^{1}\frac{ 2(y+1)^{3/2} }{ 3 }\] \[\frac{ 2 }{ 3 }(1+1)^{3/2}-\frac{ 2 }{ 3 }(0+1)^{3/2} \]Which comes out to be something with two terms and a radical in the numerator in the first, it doesn't really matter, but it's definitely incoorect; the answer is 2/3.

OpenStudy (anonymous):

all good till the very last step.. \[ \int_0^1u^{1/2}du={2\over3}[u^{3/2}]_0^1={2\over3}(1-0)=\mathbf{2\over3} \]

OpenStudy (anonymous):

you do not convert it back to "y".

OpenStudy (anonymous):

but if you DO convert it, then your limits change back to the original limits

OpenStudy (mendicant_bias):

Just solved this during lunch with a friend, thanks, though. Well, you can convert it, right? It's just a choice of whether to do that or not, even though not converting it is much less tedious and much more sensible.